A football is punted from a height of 2.5 feet above the ground with an initial vertical velocity of 45 feet per second.
1.) Write an equation to model the height h in feet of the ball t seconds after it has been punted.
2.) The football is caught at 5.5 feet above the ground. How long was the football in the air?

Answer by ankor@dixie-net.com(22740)

A football is punted from a height of 2.5 feet above the ground with an initial vertical velocity of 45 feet per second.
:
1.) Write an equation to model the height h in feet of the ball t seconds after it has been punted.
h = -16t^2 + 45t + 2.5
:
2.) The football is caught at 5.5 feet above the ground. How long was the football in the air?
-16t^2 + 45t + 2.5 = 5.5
-16t^2 + 45 t + 2.5 - 5.5 = 0
-16t^2 + 45t - 3 = 0
solving this using the quadratic formula; a=-16; b=45; c=-3, a got a reasonable solution of
t = 2.744 sec